Suppose 9.00 mol of hydrogen reacted with sufficient nitrogen. How many moles of ammonia would be produced?


1
Expert's answer
2021-03-11T08:46:34-0500

Solution:

The balanced chemical equation:

3H2 + N2 → 2NH3

According to the equation above: n(H2)/3 = n(NH3)/2

Hence,

n(NH3) = 2 × n(H2) / 3

n(NH3) = (2 × 9.00 mol) / 3 = 6.00 mol

n(NH3) = 6.00 mol


Answer: 6.00 mol of ammonia (NH3) would be produced.